d. The complete mechanism of the reaction is shown below (actually, Ive simplified it just a bit). Hydrazine (N2Ha) is an
Answer #1

d. This reaction is an Exothermic Process i.e. heat is released during the reaction.

Overall reaction can be represented as

  N_{2}\, + \, 3H_{2}\, \rightleftharpoons \, 2NH_{3}\, \, \, \, \, \Delta \,H^{o}\, =\, -91.8\, kJ/mol

Therefore reaction coordinate diagram for this is equation is:

Energy N2+ 3H2 AHE-91.8 ku/mol Reaction progress

e. Instantaneous law can be defined as change reactants concentration or of products at any instant time in the reaction.

Instant rate can be calculated using  

3 Instantaneous reaction rate dt

where dA = small change in concentration of a NH3 (reactant or product)

dt = small change in time at that instant

f. It can be seen from above equation that for 2g of H2 reacts with 14g of N2 to produce 51g of NH3

so comparing with given data for

100g of N2 = (2/14)*100 of H2 is required = 14.28g of H2

therefore N2 is the limiting reactant

Theoretical yield of above reaction is,

100g of N2 = (51/14)*100 g of NH3 = 364.28g of NH3

%yield = (Actual yield/ Theoretical yield)*100

= (45/364.28)*100 = 12.33%

g. i.  On increasing the external pressure according to Le Chatelier's principle the equilibrium will move towards the right resulting in higher yield of NH3. Pressure change cause system to adjusts in order to reduce the effect of the change since more number of molecules cause high pressure.

ii. Increase in reaction temperature generally favours endothermic reaction in order to cool system to required temperature therefore this will result in shift of equilibrium towards left side i.e production of NH3 will decrease.

iii On addition of catalyst in reaction the forward and reverse reaction rate will gradually increase but the concentrations of system remains same therefore equilibrium remains unaffected because of catalyst since catalyst only speeds up the reaction of the system by lowering the activation energy.

iv Reaction will be shift forward i.e.towards right in order to attain equilibrium with all three components, by producing NH3 molecules all over again.

h. Equilibrium expression of this reaction is

Kc = [NH3(g)]2     
[N2(g)][H2(g)]3

hence using this equation

0.453 = [NH3]2/ 2

NH3 = 0.951 M